# Harnessing Vision-Language Models for Time Series Anomaly Detection

### v1

@@ Abstract
Time-series anomaly detection (TSAD) has played a vital role in a variety of fields, including healthcare, finance, and industrial monitoring. Prior methods, which mainly focus on training domain-specific models on numerical data, lack the visual–temporal reasoning capacity that human experts have to identify contextual anomalies. To fill this gap, we explore a solution based on vision language models (VLMs). Recent studies have shown the ability of VLMs for visual reasoning tasks, yet their direct application to time series has fallen short on both accuracy and efficiency. To harness the power of VLMs for TSAD, we propose a two-stage solution, with (1) ViT4TS, a vision-screening stage built on a relatively lightweight pre-trained vision encoder, which leverages 2-D time series representations to accurately localize candidate anomalies; (2) VLM4TS, a VLM-based stage that integrates global temporal context and VLM’s reasoning capacity to refine the detection upon the candidates provided by ViT4TS. We show that without any time-series training, VLM4TS outperforms time-series pre-trained and from-scratch baselines in most cases, yielding a 24.6% improvement in F1-max score over the best baseline. Moreover, VLM4TS also consistently outperforms existing language model-based TSAD methods and is on average 36$\times$ more efficient in token usage.
1 Introduction
Time series anomaly detection (TSAD) is an important task for maintaining safety and efficiency in many domains, such as cloud computing, industrial monitoring, and web services [1]. One critical challenge is that time series signals usually exhibit diverse temporal scales and dynamic behaviors, demanding deep temporal understanding to distinguish true anomalies from benign fluctuations. For example, a sudden increase in spacecraft telemetry readings may be benign if it has recurred frequently in historical records, whereas a gradual drift that deviates from an established trend could be an anomaly [2]. However, most existing TSAD models are built based on domain-specific assumptions and are trained on numerical data, limiting them to detecting rather surface-level anomalies and cannot provide rich visual-temporal reasoning capacities that human experts have [3].
@@ 1 Introduction
Recent breakthroughs in multimodal foundation models have demonstrated human-level reasoning across multiple modalities such as images, audio, and video [4]. However, there is not yet a well-developed time-series-language model that brings human-like reasoning capacities for time series analysis and anomaly detection, largely because time series–text corpora are scarce [3]. Vision–language models (VLMs) have become a good alternative for two main reasons. First, their vision transformer architecture mimics a human-expert visual inspection mechanism—comparing and contrasting patterns at multiple scales to localize and characterize anomalies. Compared with LLM-based approaches [5], such a vision-based mechanism is especially effective at examining contextual anomalies that don’t include extreme values but show deviation from normal signal patterns, see Figure 1-(a). Second, pretraining on massive image–text datasets equips VLMs with strong inference and reasoning capabilities that generalize across domains [6]. By encoding time series as visual representations, we can leverage VLMs’ rich visual–textual understanding to perform TSAD tasks.
Figure 2: (Upper) max F1 score averaged over all benchmarks, comparing VLM4TS to time-series-pretrained and from-scratch baselines. Error bars indicate standard deviation across datasets. (Lower) max F1 score and token usage of VLM4TS versus language-model-based baselines.
@@
Recently, a few pioneering works have explored VLM-based TSAD by rendering raw time series as line graphs with x-axis tick marks, then prompting the model to return anomaly intervals by referencing those tick labels [7, 8, 9]. In practice, however, such a naive application of VLMs faces a major resolution-context dilemma when selecting the window size, as demonstrated in Figure 1-(b). With short context windows, we face challenges with high cost, latency, and limited context. To not only keep tick marks legible but also satisfy VLM input constraints (e.g. GPT-4o’s 2048-pixel maximum), long series must be split into many small rolling window plots. For example, a 1,000-step sequence might produce around 20 images, consuming around 20,000 tokens and incurring high latency, which is untenable for real time monitoring across thousands of sensors [10]. Furthermore, each plot captures only a brief temporal segment, thus preventing reasoning over global trends or long range dependencies that distinguish true anomalies from transient noise. On the other hand, with long context windows, we face challenges with limited resolution, poor localization, and missed anomalies. Plotting long series as one single image reduces token usage but significantly degrades the image resolution. The VLM cannot pinpoint the exact interval boundaries via crowded x-axis ticks, even if an anomalous behavior is identified. Moreover, the volume of visual content overwhelms the attention of the model, causing it to overlook some of the anomalies (See Figure 5 for more examples).
@@
In this paper, our goal is to harness the power of VLMs for TSAD by addressing the aforementioned resolution–context dilemma in a cost-effective manner. We propose a novel framework to decompose the anomaly detection task into two components: localization and verification. The first step localizes anomalies via a lightweight pretrained vision encoder (ViT4TS), while the second step verifies anomalies with a heavier yet more powerful VLM (VLM4TS). Specifically, for the first stage, ViT4TS adapts a pretrained vision transformer to sliding-window plots of the series, using multi-scale cross-patch comparisons to generate anomaly candidates with accurate localization. In the second stage, VLM4TS takes the candidate proposals, renders a larger plot, and prompts a large VLM to verify the anomalies and refine the detection by leveraging its multi-modal reasoning capacity. We emphasize that both stages of our framework work in a zero-shot setting—that is, neither ViT4TS nor VLM4TS is fine-tuned on the in-domain time series data. ViT4TS uses off-the-shelf, image-pretrained weights to screen windows purely via visual pattern matching, and VLM4TS relies on the pretrained cross-modal reasoning of a VLM, showing great generalizability to a broad range of TSAD tasks. Figure 2 showcases the gain in detection performance and token efficiency of our approach. To our knowledge, this is the first VLM-based TSAD solution to achieve both superior detection accuracy and practical computational and token efficiency. To summarize, our main contributions are listed as follows.
@@
$\bullet$ Motivated by human diagnostic workflows, we explore a solution that casts a 1-D anomaly detection as a 2-D visual reasoning problem with VLMs. To resolve the resolution–context dilemma in VLM application, we explore decomposing the problem into sequential localization and verification stages.
@@
$\bullet$ For the localization task, we propose ViT4TS, which leverages rich multi-scale cross-patch comparison to accurately localize anomaly candidates; then, for the verification task, we propose VLM4TS, that produce a final detection by filtering and refining on the detected anomalous intervals raised by ViT4TS with a deep temporal reasoning on global temporal context.
@@ 2 Related Work
In-domain TSAD. Early in-domain TSAD methods relied on statistical and distance-based techniques—e.g., threshold detectors and ARIMA residual analysis [11], which demand clean, uncontaminated training data and extensive domain expertise. Later on, many deep learning-based methods were introduced, applying structures like RNNs and sequence autoencoders to detect anomalies via prediction or reconstruction error [2, 12, 13]. Some works also explored VAE- and GAN-based frameworks to model normal behavior probabilistically [14, 15, 16, 17]. More recently, Transformer-based models have shown the capacity to capture more complex temporal patterns for detection [18, 19]. For a comprehensive survey, see [20]. Although highly effective in-domain, these approaches require large training sets and often fail without sufficient training and may not generalize well beyond the conditions they were trained on.
@@
Foundation Model-based and LLM-based TSAD. Recent efforts pretrain general-purpose time series encoders [21, 22] or build large forecasting foundation models [23, 24, 25, 26, 27] that can be adapted for anomaly detection. However, these models usually excel at forecasting the next few time steps or encoding a short time window, but lack the capacity to perform human-like reasoning over long-range temporal context, making them inefficient for detecting contextual anomalies across extended horizons. Prompting LLMs directly on numerical sequences has also been explored [5], yet they have shown that naive prompting underperforms some of the classical TSAD methods and comes with huge token costs.
@@
Vision and VLM-based TSAD. Treating time series as images has shown promise in classification [28, 29, 30] and forecasting [31, 32, 33], but the application in unsupervised anomaly detection remains relatively underexplored. There are few works exploring vision-based TSAD approaches using spectrograms or RP transformations [34, 35], yet these pure-vision-based approach does not and can hardly extend to leverage the cross-modal and temporal reasoning capacities of VLMs. Recent prototypes like TAMA [7] and related methods [8, 9] prompt VLMs on small rolling-window plots to assign anomaly scores, but they incur prohibitive token costs and offer only limited temporal context, as discussed in Section 1. Consequently, they cannot scale to real-world monitoring tasks or capture contextual anomalies based on long-range temporal reasoning required for robust anomaly detection.
3 Methodology
@@
To overcome the resolution–context dilemma and token inefficiency of naive VLM-based TSAD, we introduce a unified, two-stage framework (Figure 3). For the first stage, we leverage a lightweight, pretrained vision transformer on sliding-window plots to rapidly screen the entire series and generate anomaly candidates via cross-patch comparisons (Section 3.1). In the second stage, we propose VLM4TS, which takes each candidate proposal, renders it at a larger temporal scale, and applies an LLM’s cross-modal understanding to verify and precisely localize anomalies across extended horizons (Section 3.2).
@@
Problem Formulation. Consider a univariate time series $\boldsymbol{x}=(x_{1},...,x_{T})\in\mathbb{R}^{T}$, where $x_{t}$ represents the value sampled at timestamp $t$. Assume there exists a set of anomalies of varied length $\mathbf{A}=\{\left(t_{s},t_{e}\right)^{i}\mid 1\leq t_{s}<t_{e}\leq T\}_{i=1}^{m}$ that is unknown a priori, our goal is to find a set of $m$ anomalous time intervals, where $t_{s}$ and $t_{e}$ represent the start and end time points of an anomalous interval. In this paper, we consider the univariate time series to focus on key method development, the framework can be extended to multivariate data—for example, by rendering a stacked or color-coded line plot that overlays multiple channels. We leave these multivariate adaptations to future work.
@@
3.1 ViT4TS: Visual Times Series Anomaly Screening
@@
Rolling Windows and Image Creation. Pretrained vision encoders usually require square inputs (e.g. $224\times 224$ or $336\times 336$ pixels), yet raw series $\mathbf{x}$ often span far more time steps than a single image width. To achieve precise, local-scale anomaly localization, we extract overlapping windows of length $L_{w}$—chosen to match the image width—using a stride $L_{s}=\lfloor L_{w}/4\rfloor$. Each segment is rendered as a clean line plot on an $L_{w}\times L_{w}$ canvas, where each time tick maps to one pixel column. To make images of different time segments directly comparable, we use the same y-axis limits set to $[\min_{t}x_{t},\max_{t}x_{t}]$ for every image. This process produces $N\approx(T-L_{w})/L_{s}+1$ images $\mathbf{I}_{i}\in\mathbb{R}^{L_{w}\times L_{w}\times 3}$ (replicated to three channels for compatibility), which are then fed into vision encoders. For brevity, we omit explicit window subscripts in $\mathbf{I}_{i}$ in later sections.
Multi-Scale Embedding Extraction. For image encoding, we adopt the pretrained CLIP image encoder $f_{\mathrm{clip}}$ [36], as it is a standard VLM vision encoder backbone [37][38] and thus can align our visual screening results with the subsequent VLM4TS stage, yet any other vision model can also be used within this framework. To capture fine-grained time series anomalies, following the standard practice in image segmentation [39, 40], for each image we extract the full patch-level penultimate feature map $\mathbf{F}=f_{\mathrm{clip}}(\mathbf{I})\;\in\;\mathbb{R}^{P\times P\times D},$ where $P$ is the patch size and $D$ is the embedding dimension. In this way, anomalies like narrow spikes or brief dips can be localized at patch resolution.
At the same time, many anomalies, such as an extended jump or a drift that distorts the temporal patterns over dozens of time steps, will span multiple patches and demand broader context. To capture features at varying spatial scales, we generate pooled feature maps $\{\mathbf{F}^{(k)}\}_{k\in\mathcal{K}}$ by applying average pooling with kernel size $k$ and stride 1 over the base patch feature map $\mathbf{F}\in\mathbb{R}^{P\times P\times D}$. Concretely, for each $k\in\mathcal{K}$ and output location $(i,j)$,
$\mathbf{F}^{(k)}_{i,j,:}=\frac{1}{k^{2}}\sum_{u=0}^{k-1}\sum_{v=0}^{k-1}\mathbf{F}_{i+u,\;j+v,\;:},\quad i,j=1,\dots,P-k+1.$
Each $\mathbf{F}^{(k)}\in\mathbb{R}^{(P-k+1)\times(P-k+1)\times D}$ aggregates overlapping $k\times k$ neighborhoods, trading fine spatial detail for broader contextual information. These scale-specific feature maps are collected for later aggregation. Another advantage of such a multi-scale embedding is to enable rich cross-window comparisons, which in turn allows a larger rolling-window stride $L_{s}$, and thus improving efficiency without sacrificing localization accuracy.
@@
$\mathbf{S}_{i}^{(k)}[p]=\mathrm{median}_{j\neq i}\,\boldsymbol{U}_{i,j}^{(k)}[p],\ \text{with}\ \ \boldsymbol{U}_{i,j}^{(k)}[p]=\min_{r}\bigl{[}1-\cos(\mathbf{L}_{i}^{(k)}[p],\mathbf{L}_{j}^{(k)}[r])\bigr{]}.$
@@
$\;\tilde{\mathbf{S}}_{i}^{(k)}[p]=\min_{r}\bigl{[}1-\cos(\mathbf{L}_{i}^{(k)}[p],\mathbf{V}^{(k)}[r])\bigr{]},\ \text{with}\ \ \mathbf{V}^{(k)}[r]=\mathrm{median}_{j}\mathbf{L}_{j}^{(k)}[r].$
@@
The median-reference variant drastically reduces memory usage, while the all-pairs variant remains more sensitive to rare anomalies. In our main experiments, we adopt the median-reference approach for its efficiency and report the min-reference results in Appendix B.1. Finally, to produce a single, fine-grained anomaly map, we upsample each scale-specific map $\{\mathbf{S}_{i}^{(k)}\}_{k\in\mathcal{K}}$ to the base patch resolution and fuse them via harmonic averaging at every patch location. This multi-scale fusion combines the pinpoint sensitivity of small-scale scores with the broader context captured at larger scales. Examples of the resulting aggregated patch-level anomaly maps are shown in Figure 3.
@@
Anomaly Scoring and Localization. To assemble a final anomaly score per time step, we first map each fused patch-level score back to its original time index and average overlapping contributions to form a 2-D anomaly map $\mathbf{M}\in\mathbb{R}^{P\times T}$, where $P$ is the patch resolution and $T$ the series length. We then collapse $\mathbf{M}$ to a 1-D score $s\in\mathbb{R}^{T}$ by taking the $q$-th quantile across all rows: $s(t)=\mathrm{quantile}_{q}\bigl{(}\mathbf{M}_{:,t}\bigr{)}.$ Lower $q$ (e.g. 0.1) increases sensitivity to subtle distortions, while higher $q$ (e.g. 0.5) emphasizes large spikes; we set $q=0.25$ for a balanced trade-off (Appendix B.1). Finally, we choose a threshold $\tau$ at the $(1-\alpha)$-Gaussian quantile of $s$ and extract all contiguous intervals $\hat{\mathbf{A}}=\{(t_{s},t_{e})^{i}\mid s(t)>\tau,\ \forall\,t_{s}\leq t\leq t_{e}\}_{i=1}^{m}$ as our detected anomalies from ViT4TS.
3.2 VLM4TS: Vision-Language Model-based Anomaly Verification and Refinement
@@
Once ViT4TS produces a set of accurate candidate intervals, we send the full series image and these candidate proposals to VLM4TS for verification under the global context. ViT4TS excels at precise, local detection but may flag benign fluctuations or miss extended anomalies; VLM4TS uses cross-modal attention to resolve these cases. In our main experiments we employ GPT-4o for its state-of-the-art multimodal reasoning capacity, and evaluate alternatives in ablation studies.
@@
Visual Input. We render each complete time series as an ordinary line graph—with evenly spaced x-axis ticks and y-axis value labels—so that the VLM can perceive trends, seasonality, and drifts at a glance. Because ViT4TS has already provided precise endpoints, we feed this single, full-length image per signal into the VLM’s vision channel, avoiding further windowing and enabling reasoning over the entire horizon.
@@
Textual Input. In parallel, we supply a prompt that (1) lists the initial proposals $\hat{\mathbf{A}}$ and remind the model these were generated from local shape matching; (2) instructs it to confirm only intervals that truly deviate from the global pattern; (3) rejects any false positives consistent with overall behavior; and (4) suggests additional intervals exhibiting clear statistical or visual irregularities that ViT4TS may have missed. We further ask the VLM to assign each interval a confidence score from 1 (low) to 3 (high). See Appendix A.4 for the full prompt.
The VLM returns a JSON object containing the refined anomaly set $\hat{\mathbf{A}}_{final}=\{(t_{s},t_{e})^{i}\}_{i=1}^{\hat{m}}$, per-interval confidence ratings, and a brief natural-language justification for each decision. We discard any interval with confidence = 1 to produce our final detection and diagnosis.
@@
Remark on Computation. By confining expensive VLM inference in a longer context rather than every sliding window, our two-stage pipeline cuts token usage by an average factor of 20$\times$–30$\times$ compared to naive rolling-window prompting (Table 2). ViT4TS (Stage 1) runs in reasonably short time with just CPU for moderate-length series (Appendix B.3), and VLM4TS (Stage 2) incurs only a few seconds of latency per series for GPT-4o—well within acceptable bounds for offline analysis or human-in-the-loop systems. In Section 4 we quantify these gains and demonstrate that we propose the first economically feasible VLM-based TSAD solution.
4 Experiments
@@
We perform an array of experiments to evaluate the performance of ViT4TS and VLM4TS on a variety of benchmarks on industrial anomaly detection. We also conduct an extensive ablation study to validate the individual effectiveness of our proposed components. Detailed experimental setups—including data preprocessing, evaluation metrics are provided in Appendix A. The full implementation is available at https://github.com/ZLHe0/VLM4TS.
4.1 Experimental Setup

### v2

@@ Abstract
Time-series anomaly detection (TSAD) has played a vital role in a variety of fields, including healthcare, finance, and sensor-based condition monitoring. Prior methods, which mainly focus on training domain-specific models on numerical data, lack the visual–temporal understanding capacity that human experts have to identify contextual anomalies. To fill this gap, we explore a solution based on vision language models (VLMs). Recent studies have shown the ability of VLMs for visual understanding tasks, yet their direct application to time series has fallen short on both accuracy and efficiency. To harness the power of VLMs for TSAD, we propose a two-stage solution, with (1) ViT4TS, a vision-screening stage built on a relatively lightweight pre-trained vision encoder, which leverages 2-D time series representations to accurately localize candidate anomalies; (2) VLM4TS, a VLM-based stage that integrates global temporal context and VLM’s visual understanding capacity to refine the detection upon the candidates provided by ViT4TS. We show that without any time-series training, VLM4TS outperforms time-series pre-trained and from-scratch baselines in most cases, yielding a 24.6% improvement in F1-max score over the best baseline. Moreover, VLM4TS also consistently outperforms existing language model-based TSAD methods and is on average 36 × more efficient in token usage.
Introduction
Time series anomaly detection (TSAD) is an important task for maintaining safety and efficiency in many domains, such as cloud computing, industrial monitoring, and web services (lavin2015evaluating). One critical challenge is that time series signals usually exhibit diverse temporal scales and dynamic behaviors, demanding deep temporal understanding to distinguish true anomalies from benign fluctuations. For example, a sudden increase in spacecraft telemetry readings may be benign if it has recurred frequently in historical records, whereas a gradual drift that deviates from an established trend could be an anomaly (hundman2018lstmdt). However, most existing TSAD models are built based on domain-specific assumptions and are trained on numerical data, limiting them to detecting surface-level anomalies without rich visual-temporal understanding capacities that human experts have (kong2025position).
Figure 1: (Left) max F1 score averaged over all benchmarks, comparing VLM4TS to the best time-series-pretrained and from-scratch baselines. Error bars indicate standard deviation across datasets. (Right) max F1 score and token usage of VLM4TS versus language-model-based baselines.
@@ Introduction
Recent breakthroughs in multimodal foundation models have demonstrated human-level understanding across multiple modalities such as images, audio, and video (wu2023multimodal). However, there is not yet a well-developed time-series-language model that brings human-like understanding capacities for time series analysis and anomaly detection, largely because time series–text corpora are scarce (kong2025position). Vision–language models (VLMs) have become a good alternative for two main reasons. First, their vision transformer architecture mimics a human-expert visual inspection mechanism—comparing and contrasting patterns at multiple scales to localize and characterize anomalies. Compared with LLM-based approaches (alnegheimish2024can), a vision-based mechanism is especially effective at examining contextual anomalies that don’t include extreme values but show deviation from normal signal patterns, see Figure 2-(a). Second, pretraining on massive image–text datasets equips VLMs with strong inference and reasoning capabilities that generalize across domains (nagar2024zero). By encoding time series as visual representations, we can leverage VLMs’ rich visual–textual understanding to perform TSAD tasks.
@@
Recently, few pioneering works have explored VLM-based TSAD by rendering raw time series as line graphs with x-axis tick marks, then prompting the model to return anomaly intervals by referencing those tick labels (zhuang2024see; zhou2024can; xu2025can). In practice, however, such a naive application of VLMs faces a major resolution-context dilemma when selecting the window size, as demonstrated in Figure 2-(b). With short context windows, we face challenges with high cost, latency, and limited context. To not only keep tick marks legible but also satisfy VLM input constraints, long series must be split into many small rolling window plots. For example, a 1,000-step sequence might produce around 20 images, consuming around 20,000 tokens and incurring high latency, which is untenable for real time monitoring across thousands of sensors (alnegheimish2025m). Furthermore, each plot captures only a brief temporal segment, thus preventing understanding over global trends or long range dependencies that distinguish true anomalies from transient noise. On the other hand, with long context windows, we face challenges with limited resolution, poor localization, and missed anomalies. Plotting long series as one single image reduces token usage but significantly degrades the image resolution. The VLM cannot pinpoint the exact interval boundaries via crowded x-axis ticks, even if an anomalous behavior is identified. Moreover, the volume of visual content overwhelms the attention of the model, causing it to overlook some of the anomalies.
@@
In this paper, our goal is to harness the power of VLMs for TSAD by addressing the aforementioned resolution–context dilemma in a cost-effective manner. We propose a novel framework to decompose the anomaly detection task into two components: localization and verification. The first step localizes anomalies via a lightweight pretrained vision encoder (ViT4TS), while the second step verifies anomalies with a heavier yet more powerful VLM (VLM4TS). We emphasize that both stages of our framework work in a zero‐shot setting—that is, neither ViT4TS nor VLM4TS is fine‐tuned on the in-domain time series data. ViT4TS uses off-the-shelf, image-pretrained weights to screen windows purely via visual pattern matching, and VLM4TS relies on the pretrained cross-modal understanding of a VLM, showing great generalizability to a broad range of TSAD tasks. Figure 1 showcases the gain in detection performance and token efficiency of our approach. To our knowledge, this is the first VLM-based TSAD solution to achieve both superior detection accuracy and practical computational and token efficiency. To summarize, our main contributions are listed as follows.
@@
$\bullet$ Motivated by human diagnostic workflows, we explore a solution that casts a 1-D anomaly detection as a 2-D visual understanding problem with VLMs. To resolve the resolution–context dilemma in VLM application, we explore decomposing the problem into sequential localization and verification stages.
@@
$\bullet$ For the localization task, we propose ViT4TS, which leverages rich multi-scale cross-patch comparison to accurately localize anomaly candidates; then, for the verification task, we propose VLM4TS, that produce a final detection by filtering and refining on the detected anomalous intervals raised by ViT4TS with a deep temporal understanding on global temporal context.
@@
In-domain TSAD. Early in-domain TSAD methods relied on statistical and distance-based techniques—e.g., threshold detectors and ARIMA residual analysis (pena2013arima), which demand clean, uncontaminated training data and extensive domain expertise. Later on, many deep learning-based methods were introduced, applying structures like RNNs and sequence autoencoders to detect anomalies via prediction or reconstruction error (hundman2018lstmdt; malhotra2016lstmae; wong2022aer). Some works also explored VAE- and GAN-based frameworks to model normal behavior probabilistically (geiger2020tadgan; li2020anomaly; kieu2019outlier; yin2020anomaly). More recently, Transformer-based models have shown the capacity to capture more complex temporal patterns for detection (xu2021anomaly; tuli2022tranad). For a comprehensive survey, see (zamanzadeh2024deep). Although highly effective in-domain, these approaches require large training sets and may not generalize well beyond the conditions they were trained on.
@@
Foundation Model-based and LLM-based TSAD. Recent efforts pretrain general-purpose time series encoders (gao2024units; zhang2024unimts) or build large forecasting foundation models (rasul2023lag; ansari2024chronos; das2024decoder; goswami2024moment; shi2024time) that can be adapted for anomaly detection. However, these models usually excel at forecasting the next few time steps or encoding a short time window, but lack the capacity to perform human-like understanding over long-range temporal context, making them inefficient for detecting contextual anomalies across extended horizons. Prompting LLMs directly on numerical sequences has also been explored (alnegheimish2024can), yet they have shown that naive prompting underperforms some of the classical TSAD methods and comes with huge token costs.
@@
Vision and VLM-based TSAD. Treating time series as images has shown promise across a range of tasks (ni2025harnessing), including classification (costa2024fusion; kaewrakmuk2024multi; li2023time) and forecasting (yang2023your; zeng2023pixels; semenoglou2023image). However, the application in unsupervised anomaly detection remains relatively underexplored. There are few works exploring vision-based TSAD approaches using spectrograms or RP transformations (namura2024training; lin2024hierarchical), yet these approaches works purely with the vision modality. As a result, these pure vision-based approaches do not possess the multimodal reasoning and temporal understanding capabilities. Recent prototypes like TAMA (zhuang2024see) and related methods (zhou2024can; xu2025can) prompt VLMs on small rolling-window plots to assign anomaly scores, but they incur prohibitive token costs and offer only limited temporal context. Consequently, they cannot scale to real-world tasks or capture contextual anomalies based on long-range temporal understanding required for robust anomaly detection.
Methodology
@@
To overcome the resolution–context dilemma and token inefficiency of naive VLM-based TSAD, we introduce a unified, two-stage framework (Figure 3). For the first stage, we propose ViT4TS, which leverages a lightweight, pretrained vision transformer on sliding-window plots to rapidly screen the entire series and generate anomaly candidates via cross-patch comparisons. In the second stage, we propose VLM4TS, which takes each candidate proposal, renders it at a larger temporal scale, and applies an LLM’s cross-modal understanding to verify and precisely localize anomalies across extended horizons.
@@
Problem Formulation. Consider a univariate time series $\boldsymbol{x}=(x_{1},...,x_{T})\in\mathbb{R}^{T}$, where $x_{t}$ represents the value sampled at timestamp $t$. Assume there exists a set of anomalies of varied length $\mathbf{A}=\{\left(t_{s},t_{e}\right)^{i}\mid 1\leq t_{s}<t_{e}\leq T\}_{i=1}^{m}$ that is unknown a priori, our goal is to find a set of $m$ anomalous time intervals, where $t_{s}$ and $t_{e}$ represent the start and end time points of an anomalous interval. In this paper, we mainly consider the univariate time series to focus on key method development; a discussion on extending the method for multivariate time series is provided in the Appendix.
@@
ViT4TS: Visual Times Series Anomaly Screening
@@
Rolling Windows and Image Creation. Pretrained vision encoders usually require square inputs (e.g. $224\times 224$ or $336\times 336$ pixels), yet raw series $\boldsymbol{x}$ often span far more time steps than a single image width. To achieve precise, local-scale anomaly localization, we extract overlapping windows of length $L_{w}$—chosen to match the image width—using a stride $L_{s}=\lfloor L_{w}/4\rfloor$. Each segment is rendered as a clean line plot on an $L_{w}\times L_{w}$ canvas, where each time tick maps to one pixel column. To make images of different time segments directly comparable, we use the same y-axis limits set to $\min_{t}x_{t},\max_{t}x_{t}]$ for every image. This process produces $N\approx(T-L_{w})/L_{s}+1$ images $\mathbf{I}_{i}\in\mathbb{R}^{L_{w}\times L_{w}\times 3}$ (replicated to three channels for compatibility), which are then fed into vision encoders. For brevity, we omit explicit window subscripts in $\mathbf{I}_{i}$ in later sections.
Multi-Scale Embedding Extraction. For image encoding, we adopt the pretrained CLIP vision encoder $f_{\mathrm{clip}}$ (schuhmann2021laion), as it is a standard VLM vision encoder backbone (NEURIPS2023_6dcf277e; Liu_2024_CVPR) and thus can align our visual screening results with the subsequent VLM4TS stage, yet any other vision model can also be used within this framework. To capture fine-grained time series anomalies, following the standard practice in image segmentation (Jeong_2023_CVPR; MishraAD), for each image we extract the full patch-level penultimate feature map $\mathbf{F}=f_{\mathrm{clip}}(\mathbf{I})\;\in\;\mathbb{R}^{P\times P\times D},$ where $P$ is the patch size and $D$ is the embedding dimension. In this way, anomalies like narrow spikes or brief dips can be localized at patch resolution.
At the same time, many anomalies, such as an extended jump or a drift that distorts the temporal patterns over dozens of time steps, will span multiple patches and demand broader context. To capture features at varying spatial scales, we generate pooled feature maps $\{\mathbf{F}^{(k)}\}_{k\in\mathcal{K}}$ by applying average pooling with kernel size $k$ and stride 1 over the base patch feature map $\mathbf{F}\in\mathbb{R}^{P\times P\times D}$. For each $k\in\mathcal{K}$ and spatial location $(i,j)$,
$\mathbf{F}^{(k)}_{i,j,d}=\frac{1}{k^{2}}\sum_{u=0}^{k-1}\sum_{v=0}^{k-1}\mathbf{F}_{i+u,\;j+v,\;d}.$
Here $i,j=1,\dots,P-k+1$ and $d=1,\dots,D$. Each $\mathbf{F}^{(k)}\in\mathbb{R}^{(P-k+1)\times(P-k+1)\times D}$ aggregates overlapping $k\times k$ neighborhoods, trading fine spatial detail for broader contextual information. These scale-specific feature maps are collected for later aggregation. Another advantage of such a multi-scale embedding is to enable rich cross-window comparisons, which in turn allows a larger rolling-window stride $L_{s}$, and thus improving efficiency without sacrificing localization accuracy.
@@
$\mathbf{S}_{i}^{(k)}[p]=\mathrm{median}_{j\neq i}\,\boldsymbol{U}_{i,j}^{(k)}[p],\ \text{with}$
$\boldsymbol{U}_{i,j}^{(k)}[p]=\min_{r}\bigl[1-\cos(\mathbf{L}_{i}^{(k)}[p],\mathbf{L}_{j}^{(k)}[r])\bigr].$
@@
$\tilde{\mathbf{S}}_{i}^{(k)}[p]=\min_{r}\bigl[1-\cos(\mathbf{L}_{i}^{(k)}[p],\mathbf{V}^{(k)}[r])\bigr],\text{with}$
$\mathbf{V}^{(k)}[r]=\mathrm{median}_{j}\mathbf{L}_{j}^{(k)}[r].$
@@
The median-reference variant drastically reduces memory usage, while the all-pairs variant remains more sensitive to rare anomalies. In our main experiments, we adopt the median-reference approach for its efficiency and report the min-reference results in the Appendix. Finally, to produce a single, fine-grained anomaly map, we upsample each scale-specific map $\{\mathbf{S}_{i}^{(k)}\}_{k\in\mathcal{K}}$ to the base patch resolution and fuse them via harmonic averaging at every patch location. This multi-scale fusion combines the pinpoint sensitivity of small-scale scores with the broader context captured at larger scales. Examples of the resulting aggregated patch-level anomaly maps are shown in Figure 3.
@@
Anomaly Scoring and Localization. To assemble a final anomaly score per time step, we first map each fused patch-level score back to its original time index and average overlapping contributions to form a 2-D anomaly map $\mathbf{M}\in\mathbb{R}^{P\times T}$, where $P$ is the patch resolution and $T$ the series length. We then collapse $\mathbf{M}$ to a 1-D score $s\in\mathbb{R}^{T}$ by taking the $q$-th quantile across all rows: $s(t)=\mathrm{quantile}_{q}\bigl(\mathbf{M}_{:,t}\bigr).$ Lower $q$ (e.g. 0.1) increases sensitivity to subtle distortions, while higher $q$ (e.g. 0.5) emphasizes large spikes; we set $q=0.25$ for a balanced trade-off (See the Appendix for details). Finally, we choose a threshold $\tau$ at the $(1-\alpha)$-Gaussian quantile of $s$ and extract all contiguous intervals $\hat{\mathbf{A}}=\{(t_{s},t_{e})^{i}\mid s(t)>\tau,\ \forall\,t_{s}\leq t\leq t_{e}\}_{i=1}^{m}$ as our detected anomalies from ViT4TS.
VLM4TS: Anomaly Verification and Refinement
@@
Once ViT4TS produces a set of accurate candidate intervals, we send the full series image and these candidate proposals to VLM4TS for verification under the global context. ViT4TS excels at precise, local detection but may flag benign fluctuations or miss extended anomalies; VLM4TS uses cross-modal understanding along with the global temporal context to resolve these cases. In our main experiments, we employ GPT-4o, and evaluate alternatives in ablation studies.
@@
Visual Input. We render each complete time series as an ordinary line graph—with evenly spaced x-axis ticks and y-axis value labels—so that the VLM can perceive trends, seasonality, and drifts at a glance. Because ViT4TS has already provided precise endpoints, we feed this single, full-length image per signal into the VLM’s vision channel, avoiding further windowing and enabling understanding over the entire horizon.
@@
Textual Input. In parallel, we supply a prompt that (1) lists the initial proposals $\hat{\mathbf{A}}$ and remind the model these were generated from local shape matching; (2) instructs it to confirm only intervals that truly deviate from the global pattern; (3) rejects any false positives consistent with overall behavior; and (4) suggests additional intervals exhibiting clear statistical or visual irregularities that ViT4TS may have missed. We further ask the VLM to assign each interval a confidence score from 1 (low) to 3 (high). See the Appendix for the full prompt. The VLM returns a JSON object containing the refined anomaly set $\hat{\mathbf{A}}_{final}=\{(t_{s},t_{e})^{i}\}_{i=1}^{\hat{m}}$, per-interval confidence ratings, and a brief natural-language justification for each decision. We discard any interval with confidence = 1 to produce our final detection and diagnosis.
@@
Remark on Computation. By confining expensive VLM inference in a longer context rather than every sliding window, our two-stage pipeline cuts token usage by an average factor of 36 $\times$ compared to naive rolling-window prompting (Table 2). ViT4TS (Stage 1) runs in reasonably short time with just CPU for moderate-length series (See the Appendix), and VLM4TS (Stage 2) incurs only a few seconds of latency per series—well within acceptable bounds for offline analysis or human-in-the-loop systems.
Experiments
@@
We perform an array of experiments to evaluate the performance of ViT4TS and VLM4TS on a variety of benchmarks on industrial anomaly detection. We also conduct an extensive ablation study to validate the individual effectiveness of our proposed components. Detailed experimental setups—including data preprocessing, evaluation metrics are provided in the Appendix.
Experimental Setup